# A theory of state-to-state transitions based on the framework of classical reaction dynamics

### v1

@@ Abstract
We propose a new method to describe the population dynamics of distinct configurational states based on a continuous-time description of state-to-state transitions. According to classical reaction dynamics theory, the probability density associated with a given state obeys the Liouville equation, including influx from and efflux to neighboring states. By introducing a Markov approximation for the crossing of boundaries separating the states, tractable integral equations governing the state populations are derived. Once the time-dependent quantities appearing in these equations are evaluated, the population dynamics on long timescales can be obtained. Because these quantities depend only on a few states in the local neighborhood of a given state, they can be computed using a set of short-timescale molecular dynamics (MD) simulations. We apply the present method to the binding and unbinding kinetics of CH4/CH4, Na+/Cl-, and 18-crown-6-ether (crown ether)/K+ in water. For both kinetics, the time constants estimated from the present method are almost comparable to those obtained from brute-force MD simulations. The required timescale of each MD trajectory in the present method is approximately two orders of magnitude shorter than that in the brute-force MD approach in the crown ether/K+ system. This reduction in the trajectory timescale enables applications to complex binding and unbinding systems whose characteristic timescales are far beyond those directly accessible by brute-force MD simulations.
@@ I Introduction
Here, we develop a methodology based on the classical dynamics theory to describe the state-to-state transitions. The integral equations of the population dynamics are derived through a systematic approximation. Conceptually, the derived equations are similar to those used in the original milestoning theory[17, 5], but they describe the populations of states rather than those on boundaries (milestones). This feature makes it possible to compute various types of time-correlation functions related to populations, such as the hydrogen-bond time-correlation function[35, 31] and first-passage time distribution functions,[66, 30] without modifying the original definitions of these functions. The time-dependent functions for a state involved in the derived equations are described only by a limited number of states in the local neighborhood, and thus their evaluation is possible using a number of short-timescale MD trajectories.
We apply the developed method to the binding and unbinding kinetics of three aqueous systems: CH4/CH4, Na+/Cl-, and 18-crown-6 ether (crown ether)/K+. Since the timescales of binding and unbinding for these systems are accessible to MD simulations, as demonstrated by Zwier et al.,[78] these systems are suitable for verifying the accuracy of kinetic quantities estimated from the present method. We perform short-timescale MD simulations to compute the quantities required for the present method. In addition, we conduct brute-force MD simulations to directly evaluate the kinetic quantities without resorting to any approximations, for comparison.
@@ II.2 Reaction model
$\hat{f}_{j}\left(\bm{\Gamma},t\right)$ $=\hat{\Psi}\left(\bm{\Gamma},t\right)\Theta_{j}\left(\bm{\zeta}\right).$ (7)
In the above expression, the value of the reaction coordinate at $\bm{\Gamma}$, $\bm{\zeta}\left(\bm{\Gamma}\right)$, is denoted simply as $\bm{\zeta}$. Evidently, Eq. (7) satisfies the following relationship:
@@
Applying the Liouville operator ($\mathcal{L}$) to $\Theta_{j}\left(\bm{\zeta}\right)$ gives
$\mathcal{L}\Theta_{j}\left(\bm{\zeta}\right)$ $=\left(\mathcal{L}\bm{\zeta}\right)\cdot\sum_{i\in\mathcal{N}_{j}}\left(-{\bf n}_{ij}\right)\delta_{ij}^{s}\left(\bm{\zeta}\right)$
$=-\sum_{i\in\mathcal{N}_{j}}\left(\dot{\bm{\zeta}}\cdot{\bf n}_{ij}\right)\delta_{ij}^{s}\left(\bm{\zeta}\right),$ (13)
where we have used the relationship given by $\mathcal{L}\bm{\zeta}=\dot{\bm{\zeta}}$. $\delta_{ij}^{s}\left(\bm{\zeta}\right)$ is the surface delta function for boundary $s_{ij}$, defined as the function satisfying
$\int d\bm{\xi}\,A\left(\bm{\xi}\right)\delta_{ij}^{s}\left(\bm{\xi}\right)$ $=\int_{s_{ij}}d\sigma\,A\left(\bm{\xi}\right),$ (14)
for an arbitrary function $A$. Here, $\int_{s_{ij}}d\sigma$ is the surface integral over $s_{ij}$. By substituting Eq. (13) into Eq. (11), one can obtain
@@ II.3 Coarse-grained reaction dynamics
$M_{ij}\left(t\right)$ $=\dfrac{\Braket{e^{-\mathcal{L}_{j}t}S_{ij}\left(\bm{\zeta},\dot{\bm{\zeta}}\right)}}{\Braket{S_{ij}\left(\bm{\zeta},\dot{\bm{\zeta}}\right)}}.$ (31)
$Q_{ij}\left(t\right)$ represents the efflux from state $j$ to state $i$ during the time interval from $t$ to $t+dt$, while $M_{ij}\left(t\right)$ denotes the population of state $i$ that arrived from state $j$ during the time interval from 0 to $dt$ and remains in state $i$ at time $t$ (Fig. 2(a)). Since $M_{ij}\left(t\right)$ is conditioned by $\Braket{S_{ij}\left(\bm{\zeta},\dot{\bm{\zeta}}\right)}$, its initial value is unity. Eq. (29) implies that the population crossing from state $k$ to state $j$ during the time interval from $\tau$ to $\tau+d\tau$ and the subsequent retention of the population up to time $t$ are described by $Q_{jk}\left(\tau\right)$ and $M_{jk}\left(t-\tau\right)$, respectively, and that the retention process proceeds independently of the crossing event. The approximation introduced in Eq. (29) is valid when each state is sufficiently wide. Substituting Eq. (29) into Eq. (26) yields
@@ II.4 Reflecting and absorbing boundary conditions
We describe how to impose various boundary conditions on the formulations of the coarse-grained reaction dynamics. As will be shown in the next section, introducing boundary conditions is useful for applying the present method to the molecular binding processes through the returning probability theory.
We first consider the reflecting boundary condition, in which the population entering given states is instantly redirected to their neighboring states. Let $\mathcal{M}_{\mathrm{reflect}}$ be the set of states that prohibit such influx. The reflecting boundary condition can then be expressed as
@@ II.5 Incorporating returning probability (RP) theory
Returning-probability (RP) theory is a rigorous framework of diffusion-influenced reactions. In this approach, molecular binding is characterized in terms of the thermodynamic and dynamic properties of the reactive states that exists during the binding process. It has proven useful to analyze various systems including host-guest binding,[26, 27] and membrane permeation.[38] In the case of binding, the reactive state is defined as the region near the barrier top on the free-energy profile separating the bound and unbound states, whereas in the case of membrane permeation, it corresponds to the membrane center. In this subsection, we describe the scheme of efficiently computing the dynamic properties required in the RP theory to evaluate the binding rate constant, $k_{\mathrm{on}}$, based on the present theory.
@@ III Computational details
Table 1: The numbers of MD trajectories and trajectory length for each run used in the present method. The values in parentheses correspond to the case of binding kinetics. When values in parentheses are not shown, the listed values are common to both binding and unbinding cases.
@@ III.2 Umbrella sampling (US)
III.3 MD simulations for the present method
In order to calculate the quantities required for the present method ($R_{ij}\left(t\right)$, $K_{ijk}\left(t\right)$, $P_{j}^{0}\left(t\right)$, and $M_{ij}\left(t\right)$), we performed a number of the short-timescale MD simulations starting from different states. To examine how the definition of states affects the binding/unbinding kinetics obtained using the present method, MD simulations corresponding to different sets of states were performed. Let $N_{\mathrm{B}}$ and $\Delta_{\mathrm{O}}$ denote the number of states belonging to the bound region (Fig. 5) and the width of the states outside the bound region (unbound states) in distance, respectively. The number of unbound states was determined such that the position of the outermost boundary separating adjacent states, $d_{\mathrm{OM}}$, satisfied $d_{\mathrm{OM}}\geq 10$$\mathrm{\AA }$. An initial configuration was prepared for each state, in which the value of $r$ lay within the range defined for that state. In the outermost state, the target molecules were placed so as to satisfy $d_{\mathrm{OM}}\leq r\leq d_{\mathrm{OM}}+2$$\mathrm{\AA }$.
@@ III.4 Brute-force MD simulations
We performed the brute-force MD simulations to evaluate the binding/unbinding kinetics for comparison with the results using the present method. The number of trajectories and trajectory length are described in Table 2.
@@ IV.1 Unbinding kinetics
Figure 6: Residence time correlation function, $P_{\mathrm{B}}\left(t\right)$, for (a) CH4/CH4, (b) Na+/Cl-, and (c) crown ether/K+. For the present method, the results obtained under the condition $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=\left(3,3~\mathrm{\AA }\right)$ are shown, where $N_{\mathrm{B}}$ and $\Delta_{\mathrm{O}}$ are the number of divisions for the bound region (Fig. 5) and the width of the states outside the bound region, respectively.
Figure 7: Statistical convergence of $\tau_{\mathrm{off}}$ with respect to the number of trajectories, for (a) CH4/CH4, (b) Na+/Cl-, and (c) crown ether/K+. For the crown ether/K+ system, the number of trajectories corresponding to the states outside the bound region (unbound states) is fixed to 200, because the value of $\tau_{\mathrm{off}}$ is found to be less sensitive to the number of trajectories in the unbound states as compared to those in the bound region.
Table 3: Residence time constants estimated from the present method and the brute-force MD simulations. For the present method, the time constants are evaluated under the condition $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=(3,3~\mathrm{\AA })$. The statistical uncertainty is provided at 95% confidence interval.
@@
We first examine the reliability of the present method in terms of the unbinding kinetics. The residence time correlation function, $P_{\mathrm{B}}\left(t\right)$, defined as
@@
Fig. 7 shows $P_{\mathrm{B}}\left(t\right)$ computed using the present method under the condition $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=\left(3,3~\mathrm{\AA }\right)$, together with the corresponding results obtained from the brute-force MD simulations. Here, $N_{\mathrm{B}}$ and $\Delta_{\mathrm{O}}$ respectively denote the number of division of the bound region and the distance width of the states outside the bound region. The time interval used for discretization, $\Delta t$ (see Sec. II.6), is set to 20 fs for the CH4/CH4 and Na+/Cl- systems and to 60 fs for the crown ether/K+ system. As shown in Fig. 7, the present method reproduces the profiles of $P_{\mathrm{B}}\left(t\right)$ obtained from the brute-force MD simulations well, although a difference is slightly discernible for the CH4/CH4 system at $t\leq 10^{-2}~\mathrm{ns}$. $P_{\mathrm{B}}\left(t\right)$ converges to zero on the timescale of $10^{-2}$$10^{-1}$ ns for both the CH4/CH4 and Na+/Cl- systems, and this timescale is comparable to that obtained from the brute-force MD simulations ($0.2$ ns). In the case of the crown ether/K+ system, on the other hand, it is noteworthy that the present method enables the estimation of the timescale on which $P_{\mathrm{B}}\left(t\right)$ converges to zero ($10^{2}$$10^{3}$ ns) from a number of short-timescale (2 ns) trajectories.
We next investigate the dependency of the unbinding kinetics obtained from the present method on the combinations of $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)$ and the number of trajectories. In this analysis, the dependency is evaluated with the unbinding time constant, $\tau_{\mathrm{off}}$, defined as
@@
The values of $\tau_{\mathrm{off}}$ under the condition $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=\left(3,3~\mathrm{\AA }\right)$ are listed in Table 3. Figure 7 illustrates how $\tau_{\mathrm{off}}$ depends on the number of trajectories, $N_{\mathrm{traj}}$, for different combinations of $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)$. $N_{\mathrm{traj}}$ denotes the number of trajectories per state used in the present method and the number of trajectories starting from the bound region in the brute-force MD simulations. In the case of the crown ether/K+ (Fig. 7(c)), the number of trajectories for the states outside the bound region is fixed to 200. For the CH4/CH4 system (Fig. 7(a)), the convergence of $\tau_{\mathrm{off}}$ is achieved when $N_{\mathrm{traj}}$ is larger than approximately 70, regardless of the choice of $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)$. It is also found that the converged value of $\tau_{\mathrm{off}}$ is almost independent of $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)$. The converged value is larger than that obtained from the brute-force MD simulations, but the deviation is below 10%. As for the Na+/Cl- system, the convergence is observed when $N_{\mathrm{traj}}\gtrsim 70$ for both the present method and the brute-force MD simulations. The $N_{\mathrm{B}}$-dependence is slightly discernible, and the converged values of $\tau_{\mathrm{off}}$ at $N_{\mathrm{B}}=1$ are virtually identical to that for the brute-force MD simulations. The quantitative improvement with decreasing $N_{\mathrm{B}}$ is a reasonable trend, as the approximation introduced in the present method (Eqs. (32) and (37)) is expected to be valid when the spacing between adjacent states is sufficiently large. In the case of the crown ether/K+ system, convergence with respect to $N_{\mathrm{traj}}$ is slower compared with the other systems, especially for $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=(3,2~\mathrm{\AA })$. Although the $\Delta_{\mathrm{O}}$-dependence is also observed, the difference in the converged values between $\left(3,3~\mathrm{\AA }\right)$ and $\left(3,4~\mathrm{\AA }\right)$ is negligibly small, and these values are satisfactorily close to that obtained from the brute-force MD simulations.

### v2

# IEPDYN: Integral-equation formalism of population dynamics

@@ Abstract
We propose the integral-equation formalism of population dynamics (IEPDYN) to describe the population dynamics of distinct configurational states. According to classical reaction dynamics theory, the probability density associated with a given state obeys the Liouville equation, including influx from and efflux to neighboring states. By introducing a Markov approximation for the crossing of boundaries separating the states, tractable integral equations governing the state populations are derived. Once the time-dependent quantities appearing in these equations are evaluated, the population dynamics on long timescales can be obtained. Because these quantities depend only on a few states in the local neighborhood of a given state, they can be computed using a set of short-timescale molecular dynamics (MD) simulations. The IEPDYN method is formulated in continuous time and therefore does not rely on a coarse-grained timescale (lag time). Consequently, kinetic quantities obtained from IEPDYN are free from lag-time dependence, which has been discussed as a limitation in other approaches. We apply the IEPDYN method to the binding and unbinding kinetics of CH4/CH4, Na+/Cl-, and 18-crown-6-ether (crown ether)/K+ in water. For both kinetics, the time constants estimated from the IEPDYN method are almost comparable to those obtained from brute-force MD simulations. The required timescale of each MD trajectory in the IEPDYN method is approximately two orders of magnitude shorter than that in the brute-force MD approach in the crown ether/K+ system. This reduction in the trajectory timescale enables applications to complex binding and unbinding systems whose characteristic timescales are far beyond those directly accessible by brute-force MD simulations.
@@ I Introduction
Here, we develop a methodology called the integral-equation formalism of population dynamics (IEPDYN), which is based on classical dynamical theory and is designed to describe state-to-state transitions. The integral equations of the population dynamics are derived through a systematic approximation. Conceptually, the derived equations are similar to those used in the original milestoning theory[17, 5], but they describe the populations of states rather than those on boundaries (milestones). This feature makes it possible to compute various types of time-correlation functions related to populations, such as the hydrogen-bond time-correlation function[35, 31] and first-passage time distribution functions,[66, 30] without modifying the original definitions of these functions. The time-dependent functions for a state involved in the derived equations are described only by a limited number of states in the local neighborhood, and thus their evaluation is possible using a number of short-timescale MD trajectories.
We apply the IEPDYN method to the binding and unbinding kinetics of three aqueous systems: CH4/CH4, Na+/Cl-, and 18-crown-6 ether (crown ether)/K+. Since the timescales of binding and unbinding for these systems are accessible to MD simulations, as demonstrated by Zwier et al.,[78] these systems are suitable for verifying the accuracy of kinetic quantities estimated from the IEPDYN method. We perform short-timescale MD simulations to compute the quantities required for the IEPDYN method. In addition, we conduct brute-force MD simulations to directly evaluate the kinetic quantities without resorting to any approximations, for comparison.
@@ II.2 Reaction model
$\hat{f}_{j}\left(\bm{\Gamma},t\right)$ $=\hat{\Psi}\left(\bm{\Gamma},t\right)\Theta_{j}\left(\bm{\zeta}\left(\bm{\Gamma}\right)\right).$ (7)
In the above expression, the value of the reaction coordinate at $\bm{\Gamma}$ is denoted by $\bm{\zeta}(\bm{\Gamma})$, and the argument $\bm{\Gamma}$ is omitted hereafter for simplicity. Evidently, Eq. (7) satisfies the following relationship:
@@
Since the time derivative of $\bm{\zeta}$ is given by $\dot{\bm{\zeta}}=\mathcal{L}\bm{\zeta}$, applying the Liouville operator ($\mathcal{L}$) to $\Theta_{j}\left(\bm{\zeta}\right)$ with the chain rule for differentiation yields
$\mathcal{L}\Theta_{j}\left(\bm{\zeta}\right)$ $=\left(\mathcal{L}\bm{\zeta}\right)\cdot\dfrac{\partial}{\partial\bm{\zeta}}\Theta_{j}\left(\bm{\zeta}\right)$
$=\dot{\bm{\zeta}}\cdot\dfrac{\partial}{\partial\bm{\zeta}}\Theta_{j}\left(\bm{\zeta}\right).$ (13)
Then, let us introduce the surface delta function for boundary $s_{ij}$, $\delta_{ij}^{s}\left(\bm{\zeta}\right)$, that satisfies
$\int d\bm{\zeta}\,A\left(\bm{\zeta}\right)\delta_{ij}^{s}\left(\bm{\zeta}\right)$ $=\int_{s_{ij}}d\sigma\,A\left(\bm{\zeta}\right),$ (14)
for an arbitrary function $A$, where $\int_{s_{ij}}d\sigma$ is the surface integral over $s_{ij}$. The derivative of $\Theta_{j}\left(\bm{\zeta}\right)$ with respect to $\bm{\zeta}$ can be expressed as
$\dfrac{\partial}{\partial\bm{\zeta}}\Theta_{j}\left(\bm{\zeta}\right)$ $=-\sum_{i\in\mathcal{N}_{j}}{\bf n}_{ij}\cdot\delta_{ij}^{s}\left(\bm{\zeta}\right).$ (15)
Note that the above equation is the generalization of the relationship between the Heaviside step function $H\left(x\right)$ and delta function $\delta\left(x\right)$ given by $dH\left(-x\right)/dx=-\delta\left(x\right)$. By substituting Eqs. (13) and (15) into Eq. (11), one can obtain
@@ II.3 Coarse-grained reaction dynamics
$M_{ij}\left(t\right)$ $=\dfrac{\Braket{e^{-\mathcal{L}_{i}t}S_{ij}\left(\bm{\zeta},\dot{\bm{\zeta}}\right)}}{\Braket{S_{ij}\left(\bm{\zeta},\dot{\bm{\zeta}}\right)}}.$ (32)
$Q_{ij}(t)$ represents the population that crosses from state $j$ to state $i$ during the time interval from $t$ to $t+dt$. while $M_{ij}\left(t\right)$ denotes the population of state $i$ that arrived from state $j$ during the time interval from 0 to $dt$ and remains in state $i$ at time $t$ (Fig. 2(a)). Since $M_{ij}\left(t\right)$ is conditioned by $\Braket{S_{ij}\left(\bm{\zeta},\dot{\bm{\zeta}}\right)}$, its initial value is unity. Eq. (30) implies that the population crossing from state $k$ to state $j$ during the time interval from $\tau$ to $\tau+d\tau$ and the subsequent retention of the population up to time $t$ are described by $Q_{jk}\left(\tau\right)$ and $M_{jk}\left(t-\tau\right)$, respectively, and that the retention process proceeds independently of the crossing event. The approximation introduced in Eq. (30) is valid when each state is sufficiently wide. Substituting Eq. (30) into Eq. (27) yields
@@ II.4 Reflecting and absorbing boundary conditions
We describe how to impose various boundary conditions on the formulations of the coarse-grained reaction dynamics. As will be shown in the next section, introducing boundary conditions is useful for applying the IEPDYN method to the molecular binding processes through the returning probability theory.
We first consider the reflecting boundary condition, in which the population entering given states is instantly redirected to their neighboring states. Let $\mathcal{M}_{\mathrm{reflect}}$ be the set of states that do not allow such influx. The reflecting boundary condition can then be expressed as
@@ II.5 Incorporating returning probability (RP) theory
Returning-probability (RP) theory is a rigorous framework of diffusion-influenced reactions. In this approach, molecular binding is characterized in terms of the thermodynamic and dynamic properties of the reactive states that exists during the binding process. It has proven useful to analyze various systems including host-guest binding,[26, 27] and membrane permeation.[38] In the case of binding, the reactive state is defined as the region near the barrier top on the free-energy profile separating the bound and unbound states, whereas in the case of membrane permeation, it corresponds to the membrane center. In this subsection, we describe the scheme of efficiently computing the dynamic properties required in the RP theory to evaluate the binding rate constant, $k_{\mathrm{on}}$, based on the IEPDYN method.
@@ III Computational details
Table 1: The numbers of MD trajectories and trajectory length for each run used in the present method (IEPDYN). The values in parentheses correspond to the case of binding kinetics. When values in parentheses are not shown, the common values are used for both binding and unbinding cases.
@@ III.2 Umbrella sampling (US)
III.3 MD simulations for the IEPDYN method
In order to calculate the quantities required for the IEPDYN method ($R_{ij}\left(t\right)$, $K_{ijk}\left(t\right)$, $P_{j}^{0}\left(t\right)$, and $M_{ij}\left(t\right)$), we performed a number of the short-timescale MD simulations starting from different states. To examine how the definition of states affects the binding/unbinding kinetics obtained using the IEPDYN method, MD simulations corresponding to different sets of states were performed. Let $N_{\mathrm{B}}$ and $\Delta_{\mathrm{O}}$ denote the number of states belonging to the bound region (Fig. 5) and the width of the states outside the bound region (unbound states) in distance, respectively. The number of unbound states was determined such that the position of the outermost boundary separating adjacent states, $d_{\mathrm{OM}}$, satisfied $d_{\mathrm{OM}}\geq 10$$\mathrm{\AA }$. An initial configuration was prepared for each state, in which the value of $r$ lay within the range defined for that state. In the outermost state, the target molecules were placed so as to satisfy $d_{\mathrm{OM}}\leq r\leq d_{\mathrm{OM}}+2$$\mathrm{\AA }$.
@@ III.4 Brute-force MD simulations
We performed the brute-force MD simulations to evaluate the binding/unbinding kinetics for comparison with the results obtained from the IEPDYN method. The number of trajectories and trajectory length are described in Table 2.
@@ IV.1 Unbinding kinetics
Figure 6: Residence time correlation function, $P_{\mathrm{B}}\left(t\right)$, for (a) CH4/CH4, (b) Na+/Cl-, and (c) crown ether/K+. For the present method (IEPDYN), the results obtained under the condition $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=\left(3,3~\mathrm{\AA }\right)$ are shown, where $N_{\mathrm{B}}$ and $\Delta_{\mathrm{O}}$ are the number of divisions for the bound region (Fig. 5) and the width of the states outside the bound region, respectively.
Figure 7: Statistical convergence of $\tau_{\mathrm{off}}$ with respect to the number of trajectories, $N_{\mathrm{traj}}$, for (a) CH4/CH4, (b) Na+/Cl-, and (c) crown ether/K+. For the crown ether/K+ system, the number of trajectories corresponding to the states outside the bound region (unbound states) is fixed to 200, because the value of $\tau_{\mathrm{off}}$ is found to be less sensitive to the number of trajectories in the unbound states as compared to those in the bound region.
Table 3: Residence time constants estimated from the present method (IEPDYN) and the brute-force MD simulations. For the IEPDYN method, the time constants are evaluated under the condition $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=(3,3~\mathrm{\AA })$. The statistical uncertainty is provided at 95% confidence interval.
@@
We first examine the reliability of the IEPDYN method in terms of the unbinding kinetics. The residence time correlation function, $P_{\mathrm{B}}\left(t\right)$, defined as
@@
Fig. 7 shows $P_{\mathrm{B}}\left(t\right)$ computed using the IEPDYN method under the condition $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=\left(3,3~\mathrm{\AA }\right)$, together with the corresponding results obtained from the brute-force MD simulations. Here, $N_{\mathrm{B}}$ and $\Delta_{\mathrm{O}}$ respectively denote the number of division of the bound region and the distance width of the states outside the bound region. As shown in Fig. 7, the IEPDYN method reproduces the profiles of $P_{\mathrm{B}}\left(t\right)$ obtained from the brute-force MD simulations well, although a difference is slightly discernible for the CH4/CH4 system at $t\leq 10^{-2}~\mathrm{ns}$. $P_{\mathrm{B}}\left(t\right)$ converges to zero on the timescale of $10^{-2}$$10^{-1}$ ns for both the CH4/CH4 and Na+/Cl- systems, and this timescale is comparable to that obtained from the brute-force MD simulations ($0.2$ ns). In the case of the crown ether/K+ system, on the other hand, it is noteworthy that the IEPDYN method enables the estimation of the timescale on which $P_{\mathrm{B}}\left(t\right)$ converges to zero ($10^{2}$$10^{3}$ ns) from a number of short-timescale (2 ns) trajectories.
We next investigate the dependency of the unbinding kinetics obtained from the IEPDYN method on the combinations of $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)$ and the number of trajectories. In this analysis, the dependency is evaluated with the unbinding time constant, $\tau_{\mathrm{off}}$, defined as
@@
The values of $\tau_{\mathrm{off}}$ under the condition $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=\left(3,3~\mathrm{\AA }\right)$ are listed in Table 3. Figure 7 illustrates how $\tau_{\mathrm{off}}$ depends on the number of trajectories, $N_{\mathrm{traj}}$, for different combinations of $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)$. $N_{\mathrm{traj}}$ denotes the number of trajectories per state used in the IEPDYN method and the number of trajectories starting from the bound region in the brute-force MD simulations. In the case of the crown ether/K+ (Fig. 7(c)), the number of trajectories for the states outside the bound region is fixed to 200. For the CH4/CH4 system (Fig. 7(a)), the convergence of $\tau_{\mathrm{off}}$ is achieved when $N_{\mathrm{traj}}$ is larger than approximately 70, regardless of the choice of $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)$. It is also found that the converged value of $\tau_{\mathrm{off}}$ is almost independent of $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)$. The converged value is larger than that obtained from the brute-force MD simulations, but the deviation is below 10%. As for the Na+/Cl- system, the convergence is observed when $N_{\mathrm{traj}}\gtrsim 70$ for both the IEPDYN method and the brute-force MD simulations. The $N_{\mathrm{B}}$-dependence is slightly discernible, and the converged values of $\tau_{\mathrm{off}}$ at $N_{\mathrm{B}}=1$ are virtually identical to that for the brute-force MD simulations. The quantitative improvement with decreasing $N_{\mathrm{B}}$ is a reasonable trend, as the approximation introduced in the IEPDYN method (Eqs. (33) and (38)) is expected to be valid when the spacing between adjacent states is sufficiently large. In the case of the crown ether/K+ system, convergence with respect to $N_{\mathrm{traj}}$ is slower compared with the other systems, especially for $\left(N_{\mathrm{B}},\Delta_{\mathrm{O}}\right)=(3,2~\mathrm{\AA })$. Although the $\Delta_{\mathrm{O}}$-dependence is also observed, the difference in the converged values between $\left(3,3~\mathrm{\AA }\right)$ and $\left(3,4~\mathrm{\AA }\right)$ is negligibly small, and these values are satisfactorily close to that obtained from the brute-force MD simulations.